# Towards Verifying Exact Conditions of Density Functional Theory Approximations

### v1

@@ III-B Verifier
The Verifier takes as input the formula $\psi$ encoding the exact condition for the DFA and a domain for the inputs to the DFA. We use the same input bounds as used in Pederson and Burke [40]; for example, for GGA functionals, the domain for $r_{s}$ is the interval $[0.0001,5]$, and that for $s$ is $[10^{-12},5]$ (to avoid division-by-zero errors). Consequently, for a GGA functional, the Verifier is trying to prove the validity of the following formula:
$\forall r_{s},s\,.\ (r_{s}\in[0.0001,5]\wedge s\in[10^{-12},5])\implies\psi.$ (17)
The validity of Equation 17 is equivalent to proving that the following formula is unsatisfiable:
$r_{s}\in[0.0001,5]\wedge s\in[10^{-12},5]\wedge\lnot\psi.$ (18)
@@
Our preliminary results showed that dReal would timeout (in 2 hours) when trying to solve the formula in Equation 18. To improve performance of Verifier, we implemented a domain-splitting technique that partitions the input domain and use dReal to solve the formula $\lnot\psi$ on each subdomain separately. This simple strategy greatly improved the performance of Verifier. Furthermore, one of the goals of XCVerifier is to determine the input regions where the DFA does not satisfy the exact condition. Thus, Verifier also performs the domain-split when dReal returns a valid input violating the given condition. This allows Verifier to isolate the regions where the DFA does not satisfy the local condition.
Algorithm 1 presents the pseudo-code for Verifier, and Figure 1(e) shows a graphical representation of output of Verifier when verifying the $E_{\rm c}$ scaling inequality condition for the LYP functional. The dReal solver is called on Line 3 to find a satisfying assignment to the formula $\varphi_{D}\wedge\psi$, where $\varphi_{D}$ is the formula encoding the domain constraints on the inputs to the DFA. If the result is UNSAT, then Verifier returns after recording that the condition was verified over domain $D$ (Line 6), indicated using   in Figure 1(e). If the result is SAT and the returned model is a valid model for $\lnot\psi$, then Verifier records that a valid counterexample was found (Line 9), else the result is inconclusive (Line 11), indicated using   and  , respectively, in Figure 1(e). It is also possible that dReal is unable to determine satisfiability within the two hour time limit. In this case, Verifier records that the verification timed out for domain $D$ (Line 13), indicated using   in Figure 1(e).
Line 14 is executed if the result is SAT or the solver timed out. Verifier calls split($D$), which partitions each input dimension of $D$ into two equal parts. The Verifier is recursively called on each subdomain $D^{\prime}$ on Line 15. We set a lower limit on the size of input domain as the base case for the recursion. On Line 2, Verifier returns if the given input domain is too small as determined by the threshold $t$. For our experiments, we set the limit for recursing after a counterexample to $0.2$ for the widths of the $r_{s}$ and $s$ ranges. For an inconclusive or timeout result, we set it to $0.625$.
@@
: XCVerifier found a counterexample showing that the DFA does not satisfy the condition.
@@ IV Experimental Results
This section evaluates the performance of XCVerifier for verifying exact conditions for DFAs, and compares the results to the prior PB approach [28].
@@ 1. RQ1
    Is XCVerifier able to verify or find counterexamples for DFAs (Section IV-B )?
@@ 2. RQ2
    How does compare XCVerifier to the PB approach (Section IV-C)?
@@ IV-A Experimental Setup
We use the following five DFAs: PBE [22], a popular non-empirical GGA DFA; SCAN [24], a fully constrained non-empirical meta-GGA DFA satisfying all known properties of DFAs; LYP [25], and empirical DFA that is a key component of several commonly-used DFAs; AM05 [32], which shows efficient and superior performance on solids [41]; and VWN RPA [33], an LDA functional.
For each of these DFAs, we consider each of the applicable conditions from Section II. Note that the Lieb-Oxford conditions only apply to functionals with both an exchange and correlation component available (e.g. PBE, SCAN).
@@ IV-B Verifying Local Conditions using XCVerifier
Table I summarizes the result of using XCVerifier to verify the local condition corresponding to each DFA exact condition in Section II for the five DFAs. Visualization of the results for PBE and LYP are shown in Figure 2 and Figure 1.
From Table I, we can see that XCVerifier rarely determines that a condition is fully satisfied () for a DFA: it determines that AM05 and VWN RPA satisfy the $E_{\rm c}$ non-positivity condition, and that VWN RPA satisfies the $E_{\rm c}$ scaling inequality condition.
For some of the DFAs, XCVerifier finds counterexamples to the conditions (). It finds counterexamples to the $T_{\text{c}}$ upper bound condition for all the DFAs except for SCAN, where it times out. It also finds counterexamples for the conjectured $T_{\text{c}}$ upper bound condition for PBE and LYP, and for all conditions but $U_{\text{c}}$ monotonicity for LYP.
[…]
Interestingly, XCVerifier also times out completely for every single condition applied to SCAN, which was designed to satisfy all known properties of DFAs.
@@
PBE  For the PBE functional, XCVerifier finds a counterexample region at $r_{s}<3.125$ and $s>1.25$ for the conjectured $T_{\text{c}}$ upper bound (Figure 2(f)). It is able to verify the entire input domain for the Lieb-Oxford bounds except for at $r_{s}<0.625$, where it times out. This is also true for the $E_{\rm c}$ non-positivity condition, which also results in some inconclusive and timeout regions. For the $E_{\rm c}$ scaling inequality, XCVerifier times out in the entire domain except for where $r_{s}>0.625$ and $s>3.75$. For the $T_{\text{c}}$ upper bound condition, it finds counterexamples in $r_{s}>0.6250875$ and times out in the rest of the region.
@@
LYP  For the $E_{\rm c}$ non-positivity condition, XCVerifier returns $s>1.25$ as the counterexample region (Figure 1(d)). For the $E_{\rm c}$ scaling inequality, it returns $r_{s}<2.19$ and $s>0.94$ (Figure 1(e)). It times out for the entire input domain for the $U_{\text{c}}$ monotonicity condition and for the $T_{\text{c}}$ upper bound condition, it verifies the part of the region where $1.8750.3125<r_{s}<$ 1.875 and $s>1.25$ and finds counterexamples in the rest of the region. For the conjectured $T_{\text{c}}$ upper bound condition, XCVerifier returns $r_{s}>0.3125$ and $s>0.9375$ as the failing region (Figure 1(f)). There are also some inconclusive and timeout regions for each of the conditions, usually bordering the counterexample regions.
@@
AM05XCVerifier verifies that AM05 satisfies $E_{\rm c}$ non-positivity for the entire region and times out completely for $U_{\text{c}}$ monotonicity. For the $E_{\rm c}$ scaling inequality, it verifies the entire input region except for $r_{s}<0.625$, where it times out. It verifies the same region for the $T_{\text{c}}$ upper bound conjectured condition and returns inconclusive for $r_{s}<0.625$. It finds no counterexamples for any of the conditions.
SCAN  For the SCAN functional, XCVerifier times out after 2 hours for all of the conditions. To check if XCVerifier is able to verify a weaker condition for SCAN, we relax the the $E_{\rm c}$ negativity condition by adding a tolerance of $0.1$ such that the verifier needs to find a model for $E_{\rm c}<0.1$ rather than $E_{\rm c}<0$. However, XCVerifier times out completely for the weakened condition as well.
@@
VWN RPAXCVerifier verifies that VWN RPA satisfies the $E_{\rm c}$ non-positivity and $E_{\rm c}$ scaling inequality for the entire input domain. For the $U_{\text{c}}$ monotonicity condition, it verifies the entire region except for where $r_{s}<0.001$, where it times out. For the $T_{\text{c}}$ upper bound, it finds counterexamples in almost the entire region except for some small subregions in the middle of the input region where it times out. For the conjectured $T_{\text{c}}$ upper bound, XCVerifier is inconclusive for $r_{s}<0.3125$. The remaining region is verified as satisfying the condition.
@@
Summary for RQ1:  In our evaluation of the seven exact conditions for the five DFAs, XCVerifier was able to verify or find counterexamples for 11 condition-DFA pairs and partially verify 7, as shown in Table I. This demonstrates the feasibility of using formal-methods techniques like XCVerifier in verifying DFA implementations. However, XCVerifier also timed out for 11 of the 29 applicable pairs, particularly for the $U_{\text{c}}$ monotonicity condition and for the SCAN functional. Thus, there is room for improvement in formal methods techniques for DFT.
@@ IV-C Comparing XCVerifier to PB
(e) $E_{\rm c}$ scaling inequality w. XCVerifier
@@
For PB: ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\mathllap{\circ}$ is a counterexample to the condition, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bullet}\mathllap{\circ}$ is a point that satisfies the condition.
@@
(b) $T_{\text{c}}$ upper bound w. PB
(c) Conj. $T_{\text{c}}$ upper bound w. PB
@@
(e) $T_{\text{c}}$ upper bound w. XCVerifier
(f) Conj. $T_{\text{c}}$ upper bound w. XCVerifier
For PB: ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\mathllap{\circ}$ is a counterexample to the condition, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bullet}\mathllap{\circ}$ is a point that satisfies the condition.
@@
Inconsistent results, $\bigotimes$, mean that the counterexample regions returned by PB and XCVerifier differ significantly. This occurs for the $T_{\text{c}}$ upper bound condition for all of the functionals except for SCAN, where XCVerifier times out. In our experiments, the inconsistencies for this condition are dramatic: the regions found by PB and XCVerifier are sometimes completely opposed. For example, for the $T_{\text{c}}$ upper bound condition for PBE (Figure 2(e)), the PB finds no counterexamples, while XCVerifier finds counterexamples in the entire region except where it times out.
The $T_{\text{c}}$ upper bound condition is the only condition that requires a limit computation, which is done numerically in PB and symbolically in XCVerifier. This may explain the discrepancy. There is also a less dramatic inconsistency for the conjectured $T_{\text{c}}$ upper bound condition for PBE Figure 2(f).
Several of the results are not inconsistent, marked with $\bigodot^{*}$, meaning that neither method finds counterexamples for the condition. This is the case for when PB finds no counterexamples and concludes that the condition is satisfied for the functional, and XCVerifier either verifies the entire region or partially times out and partially verifies the region. In general, this tends to looks like XCVerifier verifying the majority of the region with a smaller timed out or inconclusive region on the border of the input domain. This is the case for the $E_{\rm c}$ non-positivity condition (e.g. for PBE Figure 2(d)), the $E_{\rm c}$ scaling inequality condition, and the Lieb-Oxford conditions when XCVerifier does not time out completely.
@@
Summary for RQ2:  Our comparison of XCVerifier to the PB approach (Table II) shows inconsistencies for 5 DFA/condition pairs, all of which were related to the two $T_{\text{c}}$ upper bound conditions. The $T_{\text{c}}$ upper bound condition requires computing limits of $\widetilde{F}_{\text{c}}$, which is a function of the potentially complicated correlation energy $\tilde{\epsilon}_{\text{xc}}$. XCVerifier computes this limit symbolically. In contrast, the PB approach uses numerical approximation, which may explain the discrepancy. This demonstrates one advantage of our approach to verifying DFT exact conditions.
@@ V Related Work
To the best of our knowledge, ours is the first work that uses formal methods to verify correctness in the context of density functional theory, an important scientific computing application. Prior work has used a testing-based approach, which are more scalable but do not provide formal guarantees. As discussed earlier, Pederson and Burke [40] use a grid-search to evaluate whether the DFA satisfies the DFT exact conditions. Lehtola and Marques [43] demonstrate the numerical instability of many commonly-used DFAs by observing slow convergence of self-consistent field (SCF) calculation with typically used numerical quadrature schemes.
@@
In general, correctness in scientific computing is recognized as a major challenge in HPC [44, 45], which needs formal methods that address the unique challenges in this domain. Progress to date includes the verification of mathematical properties in a conjugate gradient solver, a finite difference stencil, and a mesh quality metric [46], PDE solvers [47], specific properties of CG [48] and LU decomposition [49], and the floating-point equivalence of manually and automatically differentiated code [50]. The above are orthogonal to our goal of verifying exact conditions of density functional theory approximations.
@@
Testing and analysis techniques for floating-point programs have been developed in the past decade. The first set of techniques are general approaches that aim to achieve high-coverage of numerical code  [51, 52], conduct differential testing of numerical libraries [53], and perform mutation testing of floating-point expressions given a real specification [54]. However, it has been shown that simply achieving high code coverage in numerical programs does not uncover numerical issues in most cases [52], and differential testing is not feasible for DFAs because most of them are unique.
@@
A large body of work focuses on performing automated error analysis of floating-point programs [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]. While some of these approaches provide sound error bounds of floating-point programs, they suffer from important limitations with respect to program size and control structures supported. In the absence of techniques that can reason about floating-point error in non-trivial programs, a rich area of research in software testing has focused on how to efficiently generate inputs that maximize error in the output of a program  [67, 52, 68], which can shed light on the potentially worst error a floating-point program could incur in. Similarly, work has proposed techniques to generate inputs that trigger floating-point exceptions [69, 70]. Calculating error bounds of DFAs is orthogonal to our goal of verifying physical and numerical properties of their implementations.
@@
Existing work has explored finding function input ranges, also referred to as regimes, with the purpose of improving the accuracy of floating-point expressions [71] or optimizing floating-point efficiency [72]. These approaches are based on either estimating error based on dynamic input sampling, or statically performing an error analysis, which have their own limitations, as described earlier.

### v5

@@ III-B Verifier
Algorithm 1 presents the pseudo-code for Verifier. The Verifier component of XCVerifier takes as input (i) the formula $\psi$ encoding the exact condition for the DFA, and (ii) a domain for the inputs to the DFA. We use the same input bounds as used in Pederson and Burke [28]; for example, for GGA functionals, the domain for $r_{s}$ is the interval $[0.0001,5]$, and that for $s$ is $[0,5]$. Consequently, for a GGA functional, the Verifier is trying to prove the validity, or satisfiability, of the following formula:
$\forall r_{s},s\,.\ (r_{s}\in[0.0001,5]\wedge s\in[0,5])\implies\psi.$ (11)
Proving the validity of Equation 11 is equivalent to proving that the following formula is unsatisfiable:
$r_{s}\in[0.0001,5]\wedge s\in[0,5]\wedge\lnot\psi.$ (12)
Verifier uses the dReal solver [30] to check the satisfiability of the above formula.
@@
Though the $\delta$-satisfiability problem is decidable in principle, in practice, the dReal solver could also timeout: it was unable to determine whether formula $\varphi$ is UNSAT or $\delta$-SAT in the given time limit. Our preliminary results showed that dReal would timeout on the formula in Equation 12 for most functional/condition pairs, even when given a time limit of 24 hours. To improve the performance of Verifier, we implement a domain-splitting technique that partitions the input domain and uses dReal to solve the formula $\lnot\psi$ on each subdomain separately. This simple strategy greatly improves the performance of Verifier. Furthermore, one of the goals of XCVerifier is to determine the input regions where the DFA implementation violates the local condition. Thus, Verifier also performs the domain-split when dReal returns a valid model: an input that indeed violates the given condition. This allows Verifier to isolate the subregions where the DFA implementation violates the local condition.
Algorithm 1 presents the pseudo-code for Verifier, and Figure 1(f) shows a graphical representation of the output of Verifier when verifying the conjectured $T_{\text{c}}$ upper bound condition for the PBE functional.
The dReal solver is called on Line 3 to find a satisfying assignment to the formula $\varphi_{D}\wedge\lnot\psi$, where $\varphi_{D}$ is the formula encoding the domain constraints on the inputs to the DFA and $\psi$ encodes the local condition for the functional. If the result is UNSAT, then Verifier returns after recording that the condition was verified over domain $D$ (Line 6). This is indicated using   in Figure 1(f). If the result is $\delta$-SAT, dReal will also return a model $x$ for the formula. In Line 8, valid(x) checks if this model is a valid counterexample by plugging the values back into $\psi$, which encodes the local condition. If the condition is indeed violated, then Verifier records it as a counterexample (Line 9), indicated using  . Occasionally, dReal may return SAT with a model that does not actually violate $\psi$. This happens due to the $\delta$-satisfiability procedure of dReal: the model must satisfy the weakened formula ($(\lnot\psi)^{\delta}$, but not necessarily the original formula $\lnot\psi$. In such cases, Verifier records the result as inconclusive (Line 11), indicated using   in Figure 1(f). We use a two hour time limit for the dReal solver; if dReal is unable to determine (un)satisfiability in this limit, Verifier interrupts it and records that the verification timed out for domain $D$ (Line 13), indicated using   in Figure 1(f). Increasing the timeout in our experiments did not enable dReal to solve more formulas.
Line 14 is executed if the result is SAT or when dReal times out. Verifier calls split($D$), which partitions each input dimension of $D$ into two equal parts. Verifier is recursively called on each subdomain $D^{\prime}$ in Line 15. We set a lower limit on the size of input domain as the base case for the recursion. On Line 2, Verifier returns if the given input domain is too small as determined by the threshold $t$; we used $t=0.05$ in our experiments.
@@
--: the condition does not apply to the DFA;
: XCVerifier found a counterexample showing that the DFA does not satisfy the local condition.
@@ IV Experimental Results
This section evaluates the performance of XCVerifier for verifying local conditions for DFA implementations in Libxc, and compares the results to the prior PB approach [28].
@@ 1. RQ1
    Is XCVerifier able to verify or find counterexamples for local conditions of DFA implementations (Section IV-B)?
@@ 2. RQ2
    How does XCVerifier compare to the PB approach (Section IV-C)?
@@ IV-A Experimental Setup
We use the following five DFAs in our experiments: PBE [22], a popular non-empirical GGA DFA; SCAN [24], a fully constrained non-empirical meta-GGA DFA satisfying all known properties of DFAs; LYP [25], and empirical DFA that is a key component of several commonly-used DFAs; AM05 [31], which shows efficient and superior performance on solids [40]; and VWN RPA [32], an LDA functional.
For each of these DFAs, we consider each of the applicable conditions from Section II. Note that the Lieb-Oxford conditions only apply to functionals with both an exchange and correlation component available (e.g., PBE, AM05 SCAN).
@@ IV-B Verifying Local Conditions using XCVerifier
Table I summarizes the result of using XCVerifier to verify the local condition corresponding to each DFA exact condition in Section II for the five DFAs. The results described below are summarized in Table I. Visualizations for LYP and PBE are shown in Figures 1 and 2, respectively.
@@
PBE  For the $E_{\rm c}$ non-positivity condition and the Lieb-Oxford bound, XCVerifier is able to verify the entire input domain except for some timed-out and inconclusive regions along the $s$-axis (Figure 1(d)). The verified region is $r_{s}>0.9375$ for the $E_{\rm c}$ non-positivity condition and $r_{s}>0.0781$ for the Lieb-Oxford bound. The Lieb-Oxford extension to $E_{xc}$ is verified for the entire input domain (Figure 1(e)). For the $E_{\rm c}$ scaling inequality, XCVerifier verifies most of the bottom third of the input domain and times out for the rest. There are also a few small inconclusive regions for high $s$. For the $T_{\text{c}}$ upper bound condition, XCVerifier verifies the bottom two-thirds of the input domain and times out for the rest.
For the conjectured $T_{\text{c}}$ upper bound, XCVerifier finds a large counterexample region covering the upper left diagonal of the input domain (Figure 1(f)). There is an inconclusive region along the border of the counterexample region. XCVerifier verifies or times out for the rest of the input domain.
Finally, XCVerifier times out everywhere for the $U_{\text{c}}$ monotonicity condition.
@@
LYP  For the $E_{\rm c}$ non-positivity condition, XCVerifier finds counterexamples at $s>1.6563$ (Figure 2(d)) and the remainder of the input domain is verified to satisfy the condition. For the $E_{\rm c}$ scaling inequality, the counterexamples are at $r_{s}<2.5$ and $s>1.4844$ (Figure 2(e)). For the $T_{\text{c}}$ upper bound condition (Figure 2(f)), the counterexamples are in a small region at $r_{s}>4.8437$ and $s>2.4219$, and for the conjectured $T_{\text{c}}$ upper bound condition, the region is $r_{s}>0.625$ and $s>1.3281$. For the $U_{\text{c}}$ monotonicity condition, XCVerifier finds counterexamples at $s>1.4844$ and $r_{s}<1.4062$. The rest of the region is partially verified or timed-out. There are some small inconclusive or timed-out regions on the borders of the counterexample regions for each of the conditions.
LYP is the only DFA where XCVerifier finds counterexamples to all applicable properties.
@@
AM05XCVerifier verifies that AM05 satisfies the $E_{\rm c}$ non-positivity and the $T_{\text{c}}$ upper bound conditions in the entire input domain. It also verifies most of the domain for the $E_{\rm c}$ scaling inequality and the conjectured $T_{\text{c}}$ upper bound conditions, but times out along the $s$-axis at $r_{s}<0.0781$ for the $E_{\rm c}$ scaling inequality and $r_{s}<0.1562$ for conjectured $T_{\text{c}}$ upper bound. For the $U_{\text{c}}$ monotonicity condition, Lieb-Oxford bound, and Lieb-Oxford extension to $E_{xc}$, XCVerifier times out for the entire input domain.
SCAN  For the SCAN functional, XCVerifier times out for all of the conditions.
@@
VWN RPAXCVerifier verifies that VWN RPA satisfies the $E_{\rm c}$ non-positivity, $E_{\rm c}$ scaling inequality, and $T_{\text{c}}$ upper bound conditions for the entire input domain. It also verifies the $U_{\text{c}}$ monotonicity condition where it timed out for several of the other functionals. For the conjectured $T_{\text{c}}$ upper bound, it verifies the entire region except along the $s$-axis at $r_{s}<0.0781$, where it returns inconclusive.
@@
Summary for RQ1:  In our evaluation of the seven exact conditions for the five DFAs, XCVerifier was able to verify or find counterexamples for 13 condition-DFA pairs and partially verify seven, as shown in Table I. This demonstrates the feasibility of using formal-methods techniques like XCVerifier in verifying DFA implementations. However, XCVerifier also timed out for 11 of the 31 applicable pairs, particularly for the SCAN functional, which was designed to satisfy all known properties of DFAs. Thus, there is room for improvement in formal-methods techniques for DFT.
@@ IV-C Comparing XCVerifier to PB
(e) Lieb-Oxford ext. w. XCVerifier
@@
For PB:   (region hatched) is a counterexample to the condition,   is a point that satisfies the condition.
@@
(b) $E_{\rm c}$ scaling inequality w. PB
(c) $T_{\text{c}}$ upper bound w. PB
@@
(e) $E_{\rm c}$ scaling inequality w. XCVerifier
(f) $T_{\text{c}}$ upper bound w. XCVerifier
For PB:   (region hatched) is a counterexample to the condition,   is a point that satisfies the condition.
@@
Several of the results are not inconsistent, marked with $\bigodot^{*}$, meaning that neither method finds counterexamples for the condition. This is the case for when PB finds no counterexamples, and XCVerifier either verifies the entire region or partially verifies and partially times out (e.g., for PBE in Figure 1(d)).
@@
Summary for RQ2:  The results of XCVerifier are consistent with the PB approach (Table II). For DFA-condition pairs for which PB finds counterexamples, XCVerifier also finds counterexamples in similar regions. For pairs where PB does not find counterexamples, XCVerifier either verifies the input domain, or partially times out and partially verifies the domain.
@@ V-A Analysis of Density Functional Theory Approximations
To the best of our knowledge, ours is the first work that uses formal methods to verify correctness in the context of density functional theory, an important scientific computing application. Prior work has used a testing-based approach, which is more scalable but does not provide formal guarantees. As discussed earlier, Pederson and Burke [28] use a grid-search to evaluate whether the DFA satisfies the DFT exact conditions. Lehtola and Marques [42] show that many recent DFAs are numerically ill-behaved by studying their accuracy in computing the total exchange-correlation energy.
V-B Correctness in High Performance Computing
@@
Correctness in scientific computing is recognized as a major challenge in HPC [43, 44], which needs formal methods that address the unique challenges in this domain. Progress to date includes the verification of mathematical properties in a conjugate gradient solver, a finite difference stencil, and a mesh quality metric [45], PDE solvers [46], specific properties of CG [47] and LU decomposition [48], and the floating-point equivalence of manually and automatically differentiated code [49]. The above are orthogonal to our goal of verifying exact conditions of density functional theory approximations. Extending our approach of using dReal to verify properties of other scientific computations is an interesting future direction.
V-C Analysis of Floating-Point Programs
@@
Many testing and analysis techniques for floating-point programs have been developed in the past decade. The first set of techniques are general approaches that aim to achieve high-coverage of numerical code  [50, 51], conduct differential testing of numerical libraries [52], and perform mutation testing of floating-point expressions given a real specification [53]. However, it has been shown that simply achieving high code coverage in numerical programs does not uncover numerical issues in most cases [51]. Additionally, differential testing is not feasible for DFA implementations because most of them are unique.
@@
A large body of work focuses on performing automated error analysis of floating-point programs [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. While some of these approaches provide sound error bounds of floating-point programs, they suffer from important limitations with respect to program size and control structures supported. In the absence of techniques that can reason about floating-point error in non-trivial programs, a rich area of research in software testing has focused on how to efficiently generate inputs that maximize error in the output of a program [66, 51, 67], which can shed light on the potentially worst error a floating-point program could incur. Similarly, work has proposed techniques to generate inputs that trigger floating-point exceptions [68, 69]. Calculating error bounds of DFAs is orthogonal to our goal of verifying physical and numerical properties of their implementations.
@@
Other existing work on floating-point programs has explored finding function input ranges, also referred to as regimes, with the purpose of improving the accuracy of floating-point expressions [70] or optimizing floating-point efficiency [71]. These approaches are based on either estimating error based on dynamic input sampling, or statically performing an error analysis, which have their own limitations, as described earlier.